OpenStudy (anonymous):

The function f is such that f(x)= 2sin^2x - 3cos^2x for 0<= x <=pi State the greatest and least values of f(x)

OpenStudy (anonymous):

sin^2x=(1-cos2x)/2 cos^2x=(1+cos2x)/2 => f(x)=(-1-5cos(2x))/2 the greatest value is when cos2x is least i.e. at cos2x=-1 f(x)=2 the least value is when cos2x is largest i.e. at cos2x=1 f(x)=-3

OpenStudy (experimentx):

find absolute maxima and absolute minima .. in the interval 1) find critical points, y'=0, check for maxima and minima 2) also check for values on boundary http://www.wolframalpha.com/input/?i=plot+2sin^2x+-+3cos^2x+from+x%3D0+to+x%3Dpi

OpenStudy (experimentx):

you know how to find critical points right??

OpenStudy (anonymous):

i dont think critical points are necessary for this kindaa problem

OpenStudy (anonymous):

@vamgadu can you tell me again, simpler, more explicit, how you solved it?

OpenStudy (anonymous):

I need to close this so, just please still help :D thanks

OpenStudy (experimentx):

it seems that you are right